# Detecting local perturbations of networks in a latent hyperbolic embedding space

### v1

@@ 3.2 Clinical data
Finally, we assessed our methods using anatomical brain connectivity networks derived from patients who underwent epilepsy surgery. We examined the brain networks of 51 patients, estimated using T1-weighted structural and diffusion-weighted MRI scans before and after surgery. The pre- and post-operative networks were obtained from an open dataset associated with the research paper by Sinha et al. [12]. The dataset included 51 individuals who underwent epilepsy surgery, from which 30 patients underwent surgery for a left temporal lobe epilepsy (TLE), and 21 for a right temporal lobe epilepsy. Within each category, participants were further divided based on post-surgical results: those with positive surgical outcomes (seizure-free after surgery) and those with negative outcomes (not seizure-free after surgery). The surgical procedure involved removing brain tissue located in the presumed epileptic zone (specifically, in the temporal lobe) to prevent seizures. In graph theory, this corresponds to cutting a subset of edges within a localized section of the network.
@@ 4.1 Synthetic networks
When the network’s connectivity is locally perturbed, we anticipate that this will be evident in a shift in the proximity relationships of the perturbed node within its embedding space, and we expect our methods to detect this change. We hypothesize that the geometric distortion will be proportional to the size of the perturbation, both locally and globally.
@@
To test this hypothesis, we apply the perturbation method described in §3.1 to selected synthetic connectivity models —specifically, small-world (Watts-Strogatz) and scale-free (Barabási-Albert) networks. To prevent particular perturbations from influencing the results, we iteratively apply single-node random perturbations a number of times proportional to the number of possible distinct perturbations. We estimate this number as $N_{p}=\binom{\bar{k}}{\lfloor\bar{k}\times m\rceil}\cdot N$, where $N$ is the network’s size, and the other factor is an estimate of the possible perturbations for a node with a degree equal to the average degree $\bar{k}$. In the binomial coefficient estimation, $m$ indicates the perturbation size as a fraction of links to cut, and $\lfloor x\rceil$ denotes the nearest integer to $x$.
We uniformly sample within the degree domain, allowing nodes with high degree $k\gg\bar{k}$ —and thus a higher number of possible perturbations— to be sampled more frequently. Additionally, we choose the sample size to be a fraction $\nicefrac{{N_{p}}}{{\alpha}}$ with $\alpha>1$, ensuring both the statistical significance of the analysis and the smoothness of the computation. For a given network model (small-world or scale-free), we also average over multiple random realizations.
In the trivial case where a perturbation of zero magnitude is applied, the embedding configuration of $G$ and $G^{\prime}$ are identical. As expected, both the kNHypD and HypErr scores yield a value of 00 uniformly across the network. Instead, the embedding configuration always changes when a non-null perturbation is applied.
@@
Figure 4: HypErr and kNHypD scores are computed for each node taken as target in a Barabási-Albert scale-free model with $N=150$ nodes and $m=8$. Same stipulations as in the caption of Fig. 3
@@
The results for the small-world network model are depicted in Fig. 3. The HypErr scores at the target nodes are significantly higher than the baseline scores, practically for all nodes’ degree classes. There is just a slight convergence of the target node’s score and the score from its second-order neighbors at the extremes of the degree domain, but there is never a reversal of the trend or an overshoot of the baseline score. The kNHypD score also correctly identifies a perturbation, except in the extremes of the nodes’ degree domain. If the targeted node has few connections, the perturbation cuts too few links relative to the original link density, and the method may not be sensitive enough. On the other hand, if the targeted node has more links, the perturbation results in a delocalization of the perturbation effect over many areas of the network, which can result in a target score comparable to the baseline.
@@
The values of the average baseline $\overline{BL}$ —defined here as the global average of the scores over all the network’s nodes— suggest that the magnitude of the perturbation definitely plays a role in the performance of the methods. As anticipated in the beginning of this section, a stronger local perturbation induces a higher distortion over larger parts of the network, which is reflected in increasing values of $\overline{BL}$. From the bottom plots in Fig. 3, we also observe that a perturbation of higher magnitude yields a larger difference between the target kNHypD score and baseline.
@@
The results for the scale-free network model are presented in Fig. 4. For HypErr score, results indicate that the localisation power becomes more stable as the perturbation magnitude increases (top plots in Fig. 4). Concerning the kNHypD score, there is still a noteworthy increase in the peak distance from the baseline, but it fails to identify the perturbation when high-degree nodes or hubs are targeted. Since this model has a power-law degree distribution, perturbing a hub leads to disturbing a larger number of distant regions of the network, decreasing the difference between the error of the perturbed node and the baseline. As shown in Fig. 4, this effect appears for both methods and is more pronounced for weak perturbations.
@@
For both network models, the scores calculated from the second-order neighbors of targeted nodes are very low and similar to those obtained from the baseline, suggesting a good localization power, as these nodes are practically unaffected by perturbations. Strong local correlations of nodes (assortativity) could, however, have an effect and induce a bias in the scores of these “second-order target nodes”. Therefore, we can conclude that: i) the proposed perturbation scores have good localization and a small bias; ii) the performance improves with perturbation magnitude; and iii) the performances of the kNHypD score depend on degree heterogeneity.
@@
Figure 5: Subplots in the first row show results from patients operated in the left hemisphere, wheres those operated in the right hemisphere are displayed in the second row. The first two columns display results from HypErr score, whereas results from kNHypD score are shown in the 3rd and 4th column. For each method, nodes were separated into the left hemisphere (left subplot) and right hemisphere (right subplot). In the different subplots, the nodes were assigned a ranking based on increasing perturbation degree $P_{deg}$ from left to right. Colour bars encode the mean score of the node across patients. The total height denotes the node’s mean pre-surgery degree, with mean post-surgery degree is indicated by a shaded gray section (always lower than the pre-surgical values).
@@ 4.2 Brain networks
[…]
As our methods aim to locate the disturbed region of a network, we anticipate a correlation of our scores with the perturbation degree distribution. Results shown in Fig. 5 confirm that nodes exhibiting the highest scores (encoded by colours that tend to be clearer), correspond to the nodes exhibiting the highest perturbation degree. When surgery affects brain regions with few connections, both perturbation scores yield very low values. These findings have been consistently observed in both hemispheres. A very important point is that nodes in the surgery hemisphere clearly display a stronger perturbation score than those located on the contralateral hemisphere, independently of the perturbation degree.
The spatial mapping of brain regions’ perturbation is illustrated in Fig 6. In both cases, one can see that nodes with the highest scores are always located in the surgery-impacted region (temporal lobe). We notice, however, that the perturbation score calculated with the HypErr method covers a large region, including all perturbed nodes; whereas the kNHypD score identifies a more reduced number of brain areas within the region directly affected by the surgery.
Figure 6: Left and top views of brain network perturbations in epileptic patients who have undergone left hemisphere surgery. Each node’s size corresponds to the surgery perturbation rate (rate of removed connections) of the node. The colour of nodes represents its score by either HypErr (method 1) or kNHypD (method 2).
@@ 5 Discussion
The characterization and localization of network perturbations have applications in various domains such as anomaly detection (in disease networks, computer networks, etc.) and brain surgery characterization. This study introduces two methods based on hyperbolic embedding to characterize and localize local network perturbations. Results from synthetic networks validate the proficiency of the proposed methods in identifying perturbed nodes. To showcase their practical utility, the methods are applied to real brain networks from epileptic patients who underwent surgery. The results reveal that both methods successfully pinpointed the surgical area as the most perturbed region. Notably, the HypErr score identified significant perturbations in a broader brain zone beyond the surgical region, whereas the kNHypD score highlighted perturbations in a more localized area.
The novelty of our methods resides in an explicit analysis in a latent hyperbolic space. We expect that the use of these low-dimensional and highly informative representations will be crucial for brain network studies beyond standard connectivity measures and Euclidean geometry. For this study, the coalescent embedding method [24] has been selected from a range of available methods [29, 30, 22]. This method encompasses various benefits, such as i) the absence of a minimization step that prevents errors introduced by local minima and also increases reproducibility (the same graph used as input will always produce the same embedding), and ii) a relatively short computational time. We notice, however, that other embedding techniques could potentially be combined with the proposed perturbation scores, and that alternative dimensional reduction methods (e.g., Laplacian Eigenmaps [20]) or hyperbolic mappings (e.g., Mercator [30], HyperMap [31], or Hydra [32], among others) might be worth considering for the representation of brain networks [33].
In conclusion, the article presents two methods that efficiently characterize network perturbations of the type link removal. Future work could explore the characterization of other types of perturbations, such as rewiring. Our approach may provide new insights in the study of perturbations of connected systems including social, technological or biological networks.
@@ Acknowledgements
Authors are grateful to Steven Rico for useful and valuable discussions during the preparation of the study.

### v2

@@ 3.2 Clinical data
Finally, we assessed our scores using anatomical brain connectivity networks derived from patients who underwent epilepsy surgery. We examined the brain networks of 51 patients, estimated using T1-weighted structural and diffusion-weighted MRI scans before and after surgery. The pre- and post-operative networks were obtained from an open dataset associated with the research paper by Sinha et al. [12]. The dataset included 51 individuals who underwent epilepsy surgery, from which 30 patients underwent surgery for a left temporal lobe epilepsy (TLE), and 21 for a right temporal lobe epilepsy. Within each category, participants were further divided based on post-surgical results: those with positive surgical outcomes (seizure-free after surgery) and those with negative outcomes (not seizure-free after surgery). The surgical procedure involved removing brain tissue located in the presumed epileptic zone (specifically, in the temporal lobe) to prevent seizures. In graph theory, this corresponds to cutting a subset of edges within a localized section of the network. The brain networks before the surgery have the following properties: a mean degree of $28\pm 2$ nodes, a standard deviation of the degree distribution of $12\pm 1$, a mean clustering coefficient of $0.57\pm 0.02$. After surgery, the perturbed networks display a mean degree of $25\pm 2$, a standard deviation of the degree distribution of $11\pm 1$, and a clustering coefficient of $0.55\pm 0.03$. Brain networks’ size is $N=114$ nodes and there are no disconnected nodes after surgery.
Figure 3: Results on synthetic networks. HypDisp and kNHypD scores are computed for each node taken as target in a Watts-Strogatz network with $N=150$ nodes, a lattice parameter of $20$ and $0.1$ rewiring probability. We visualize the average values per degree class, measured at the target node itself (in orange), and at the level of its nearest topological neighbors (in green), both relative to the global average over the network, the baseline (dashed grey line). In each subplot $\overline{BL}$ indicates the total average of the scores over all the network’s nodes. The bandwidths indicate the $95\%$ confidence interval of the mean. For the estimation of kNHypD, we set the number of neighbors equal to 10.
@@ 4.1 Synthetic networks
When the network’s connectivity is locally disrupted, we anticipate that this will be apparent in a change in the proximity relationships of the disrupted node within its embedding space, and we expect our scores to detect this change. We hypothesize that the geometric distortion will increase statistically with the size of the disruption, both locally and globally.
@@
To test this hypothesis, we apply the perturbation method described in §3.1 to selected synthetic connectivity models —specifically, small-world (Watts-Strogatz) and scale-free (Barabási-Albert) networks. To avoid specific perturbations from influencing the results, we repeatedly apply random perturbations to a single node a number of times that is proportional to the number of possible distinct perturbations. We estimate this number as $N_{p}=\binom{\bar{k}}{\lfloor\bar{k}\times m\rceil}\cdot N$, where $N$ is the network’s size, and the other factor is an estimate of the possible perturbations for a node with a degree equal to the average degree $\bar{k}$. In the binomial coefficient estimation, $m$ indicates the perturbation size as a fraction of links to cut, and $\lfloor x\rceil$ denotes the nearest integer to $x$.
@@
Figure 4: HypDisp and kNHypD scores are computed for each node taken as target in a Barabási-Albert model with $N=150$ nodes and $m=8$. Same stipulations as in the caption of Figure 3
We uniformly sample within the degree domain, allowing nodes with high degree $k\gg\bar{k}$ —and thus a higher number of possible perturbations— to be sampled more frequently. Additionally, we choose the sample size to be a fraction $\nicefrac{{N_{p}}}{{\alpha}}$ with $\alpha>1$, ensuring both the statistical significance of the analysis and the smoothness of the computation. For a given network model (small-world or scale-free), we also average over $n_{G}$ random network realizations.
In the trivial case where a perturbation of zero magnitude is applied, the embedding configuration of $G$ and $G^{\prime}$ are identical. As expected, both the kNHypD and HypDisp scores yield a value of 00 uniformly across the network. Instead, the embedding configuration always changes when a non-null perturbation is applied.
To quantify how much the embedding configuration is distorted locally and globally, for each perturbation we compute: i) the two geometric scores at the target node; ii) the average value of the scores relative to the nearest neighbors of the target node, i.e. the nodes sharing an edge with the target node (they are just after the target node, the most physically affected by the perturbation). iii) the baseline (BL) defined as the average of the perturbation scores across the entire network. We then divide the results obtained for the $n_{p}$ perturbations over $n_{G}$ random network realizations based on the degree of the perturbed node. This enables us to observe the perturbation as a function of degree. Finally, for each degree class, we compute the average values of (i) with respect to the baseline (iii), indicated in the figure as “Target score avg - BL”. Similarly, we obtain the average value of (ii) with respect to BL, referred as “Target neighbors score avg - BL”. The measure $\overline{BL}$ reported in each frame indicates the value of the baseline further averaged over all the degree classes.
@@
The results for the small-world network model are shown in Figure 3. The HypDisp scores at the target nodes are significantly higher than the baseline scores, practically for all nodes’ degree classes. There is only a slight convergence of the target node’s score and the score from its second-order neighbors at the extremes of the degree domain, but there is never a reversal of the trend or an overshoot of the baseline score. The kNHypD score also correctly identifies a perturbation, except in the extremes of the nodes’ degree domain. If the targeted node has few connections, the perturbation cuts too few links relative to the original link density, and the score may not be sensitive enough. On the other hand, if the targeted node has more links, the perturbation results in a delocalization of the perturbation effect over many areas of the network, which can result in a target score comparable to the baseline.
@@
The values of the average baseline $\overline{BL}$ —defined here as the global average of the scores over all the network’s nodes and over all network realizations and perturbations— suggest that the magnitude of the perturbation definitely plays a role in the performance of the scores. As anticipated in the beginning of this section, a stronger local perturbation induces a higher distortion over larger parts of the network, which is reflected in increasing values of $\overline{BL}$ (this effect is reported in Section 2 of the Supplementary Material). From the bottom plots in Figure 3, we also observe that a perturbation of higher magnitude yields a larger difference between the target kNHypD score and baseline.
@@
The results for the scale-free network model are presented in Figure 4. For HypDisp score, results indicate that the localization power becomes more stable as the perturbation magnitude increases (top plots in Figure 4). Regarding the kNHypD score, there is still a notable increase in the peak distance from the baseline, but it fails to identify the perturbation when high-degree nodes or hubs are targeted. Since this model has a power-law degree distribution, perturbing a hub leads to disturbing a larger number of distant regions of the network, decreasing the difference between the error of the perturbed node and the baseline. As shown in Figure 4, this effect appears for both scores and is more pronounced for weak perturbations.
@@
For both network models, the scores calculated from the second-order neighbors of targeted nodes are very low and similar to those obtained from the baseline, suggesting a good localization power, as these nodes are practically unaffected by perturbations. Strong local correlations of nodes (assortativity), however, could have an effect and induce a bias in the scores of these “second-order target nodes”. Interestingly, at the average degree kNHypD scores show a peak in Watts-Strogatz networks and a crossing with the baseline in the Barabási-Albert networks. In contrast, no clear trend was found for the kNHypD score as a function of the target node’s clustering coefficient (results in the Supplementary Material). Therefore, we can conclude that: i) the proposed perturbation scores have good localization and a small bias; ii) the performance improves with perturbation magnitude; and iii) the performances of the kNHypD score depend on degree heterogeneity.
@@
Figure 5: Subplots in the first row show results from patients operated in the left hemisphere, whereas those operated in the right hemisphere are displayed in the second row. The first two columns display results from HypDisp score, whereas results from kNHypD score are shown in the 3rd and 4th columns. For each score, nodes were separated into the left hemisphere (left subplot) and right hemisphere (right subplot). In the different subplots, the nodes were assigned a ranking based on increasing perturbation degree $P_{deg}$ from left to right. Color bars encode the mean score of the node across patients. The total height denotes the node’s mean pre-surgery degree, with mean post-surgery degree is indicated by a shaded gray section (always lower than the pre-surgical values).
In the calculation of the kNHypD score we have taken into account the displacement of the 10 nearest geometric neighbors of each node, i.e. $k=10$. In the Supplementary Material, you can find the results for two extreme values of this parameter ($k=1$ and $k=100$).
@@ 4.2 Brain networks
To illustrate the effectiveness of our scores on real data and compare their performances, we used them on brain network connectivity perturbations due to epilepsy surgery (refer to §3.2 for further details). For illustration purposes, here we considered patients with a positive surgical outcome. For these patients, we compared two networks: the pre- and post-surgery networks. The network after surgery is a modified version of the original network, with fewer edges specifically removed in the surgery area. A way to assess the perturbation of node $i$ due to surgery is the perturbation degree, $P_{deg}(i)$, defined as $P_{deg}(i)=\nicefrac{{(D_{pre}(i)-D_{post}(i))}}{{D_{pre}(i)}}$, where $D_{pre}(i)$ and $D_{post}(i)$ denote the pre- and post-surgery degree of the ithnode. When $P_{deg}(i)\rightarrow 0$, the node’s connectivity is slightly affected by the surgery. In contrast, when $P_{deg}(i)\rightarrow 1$, the node is more likely to become disconnected and significantly impacted.
As our scores aim to locate the disturbed region of a network, we anticipate a correlation of our scores with the perturbation degree distribution. Results shown in Figure 5 confirm that nodes exhibiting the highest scores (encoded by colors that tend to be clearer), correspond to the nodes exhibiting the highest perturbation degree. When surgery affects brain regions with few connections, both perturbation scores yield very low values. These findings have been consistently observed in both hemispheres. A very important point is that nodes in the surgical hemisphere clearly display a stronger perturbation score than those located on the contralateral hemisphere, independently of the perturbation degree.
The spatial mapping of brain regions’ perturbation is illustrated in Figure 6. In both cases, one can see that nodes with the highest scores are always located in the surgically impacted region (temporal lobe). We notice, however, that the perturbation score calculated with the HypDisp score covers a large region, including all perturbed nodes; whereas the kNHypD score identifies a more reduced number of brain areas within the region directly affected by the surgery.
Figure 6: Left and top views of brain network perturbations in epileptic patients who have undergone left hemisphere surgery. Each node’s size corresponds to the surgery perturbation rate (rate of removed connections) of the node. The color of nodes represents its score by either HypDisp (score 1) or kNHypD (score 2).
@@ 5 Discussion
The characterization and localization of network perturbations have applications in various domains such as anomaly detection (in disease networks, computer networks, etc.) and brain surgery characterization. This study introduces two scores based on hyperbolic embedding to characterize and localize local network perturbations. Results from synthetic networks validate the proficiency of the proposed scores in identifying perturbed nodes. To showcase their practical utility, the scores are applied to real brain networks from epileptic patients who underwent surgery. The results reveal that both scores successfully pinpointed the surgical area as the most perturbed region. Notably, the HypDisp score identified significant perturbations in a broader brain zone beyond the surgical region, whereas the kNHypD score highlighted perturbations in a more localized area.
The novelty of our scores resides in an explicit analysis in a latent hyperbolic space. We expect that the use of these low-dimensional and highly informative representations will be crucial for brain network studies beyond standard connectivity measures and Euclidean geometry.
For this study, the coalescent embedding method [24] has been selected from a range of available methods [25, 22]. This method encompasses various benefits, such as i) the absence of a minimization step that prevents errors introduced by local minima and also increases reproducibility (the same graph used as input will always produce the same embedding), and ii) a shorter computational time with respect to methods involving a likelihood maximization like Mercator [26]. The latter, based on a generative model, has the advantage of producing more meaningful representations. This could allow, for instance, to generate synthetic networks with similar properties to the input network. However, for the purposes of this paper, we have decided to prioritize speed and versatility over this feature. A comparison of the computational time of embeddings using coalescent and Mercator methods can be found in the Supplementary Material.
Similarly, we notice that other embedding techniques could potentially be combined with the proposed perturbation scores, and that alternative dimensional reduction methods (e.g., Laplacian Eigenmaps [17]) or hyperbolic mappings (e.g., Mercator [26], HyperMap [31], or Hydra [32], among others) might be worth considering for the representation of brain networks [33].
In conclusion, the article presents two scores that efficiently characterize network perturbations of the type link removal. Future work could explore the characterization of other types of perturbations, such as rewiring. Our approach may provide new insights in the study of perturbations of connected systems including social, technological or biological networks.
Supplementary Material
See the supplementary material for: (1) the dependency of the kNHypD score on the value of parameter $k$, (2) the evolution of the mean baseline ($\overline{BL}$) with the perturbation degree, (3) a comparison of computation times, as a function of network size, of the Mercator and the coalescent hyperbolic embedding method, and (4) the dependency of the kNHypD score on the clustering coefficient.
@@ Acknowledgements
Authors are grateful to Steven Rico for useful and valuable discussions during the preparation of the study. A.L. was supported by the doctoral school EDITE from Sorbonne University. M.G. acknowledges doctoral support from the Ecole Normale Supérieure. We thank the two anonymous referees for helpful comments.